need to solve 2x-(-5)-(-3)=x+10
Answer by stanbon(75887)

need to solve 2x-(-5)-(-3)=x+10
---
2x+5+3 = x+10
---
2x+8 = x + 10
Subtract x from both sides:
Subtract 8 from both sides:
x = 2
===================
Cheers,
Stan H.